OpenStudy (precal):

A ladder 25 feet long is leaning against the wall of a house. The base of the ladder is pulled away from the wall at a rate of 2 feet per second.

sammixboo (sammixboo):

Is this all the info?

OpenStudy (anonymous):

You may want to draw a picture.

OpenStudy (precal):

Part A. How fast is the top of the ladder moving down the wall when the base of the ladder is 7 feet from the way?

OpenStudy (anonymous):

However, I am guessing you use : \[ z^2=x^2+y^2 \]

OpenStudy (precal):

I have 2 other parts. I have solve part A. \[\frac{ db }{ dt}=-12.4375 \]

OpenStudy (precal):

feet per second

OpenStudy (precal):

Part B: Consider the triangle formed by the side of the house, the ladder, and the ground. Find the rate at which the area of the triangle is changing when the base of the ladder is 7 feet from the wall.

OpenStudy (precal):

I used \[A=\frac{ 1 }{ 2}bh\]

OpenStudy (precal):

\[\frac{ dA }{ dt }=\frac{ 1 }{ 2 }h \frac{ db }{ dt}+\frac{ 1 }{ 2 }b\frac{ dh }{ dt}\]

OpenStudy (precal):

\[\frac{ dA }{ dt }=-142.25\]

OpenStudy (precal):

feet^2/second not sure this solution is correct

OpenStudy (anonymous):

You have b, h, b', h', right? so what is the issue?

OpenStudy (precal):

I am not sure I did this one correctly nor the last part which involves the angle changing

OpenStudy (anonymous):

Which part are you unsure about?

OpenStudy (precal):

the last part the solution and the units I tend to get the units incorrect

OpenStudy (anonymous):

area/time is feet^2/sec

OpenStudy (precal):

thanks did I get the solution correct as well?

OpenStudy (precal):

I know I needed to do part A first because I needed that solution to figure out Part B

OpenStudy (precal):

Part C I am really stuck on Find the rate at which the angle formed by the ladder and the wall of the house is changing when the base of the ladder is 9 feet from the wall

OpenStudy (anonymous):

Hmm, what did you get for h, b, h'?

OpenStudy (precal):

h=24 b=7 h'=2 because that was given and b'=-12.4375 which I solved for in part A

OpenStudy (anonymous):

The base is the opposite side to the angle formed by the wall and latter. \[ 25\sin(\theta) = b \]

OpenStudy (precal):

ok so now we are on to the last part, correct. I am sorry I was still answering part b but that is fine, I can move on to part C

OpenStudy (precal):

is this the angle I am looking for?

OpenStudy (anonymous):

I would think so, since it is formed between wall and ladder.

OpenStudy (precal):

Can I use sine , cosine or tangent?

OpenStudy (precal):

\[\sin \theta=\frac{ x }{ 25 }\]

OpenStudy (anonymous):

Yes

OpenStudy (precal):

I was going to take the derivative of it

OpenStudy (precal):

\[\cos \theta \frac{ d \theta }{ dt}=\frac{ 1 }{ 25}\frac{ dx }{ dt}\]

OpenStudy (anonymous):

\[ \cos\theta = \frac y{25} \]

OpenStudy (precal):

\[\cos \theta=\frac{ \sqrt{544} }{ 25}\]

OpenStudy (precal):

I can sub for cosine theta, but I don't know dx/dt unless I use the same procedure for part A

OpenStudy (schrodingers_cat):

Not to be the bearer of bad news you need to recalc your answer for part a.

OpenStudy (precal):

Is part A incorrect?

OpenStudy (anonymous):

I never saw your work for A, and I didn't check it either. What did you do?

OpenStudy (schrodingers_cat):

625 = x^2 +y^2 0 = 2xx' + 2yy' y' = -xx'/y = -14/24

OpenStudy (precal):

yes d/dx of 625 is 0

OpenStudy (schrodingers_cat):

d/dt of constant = 0

OpenStudy (precal):

that will change part B as well

OpenStudy (precal):

Thanks part A is -7/12 ft/sec

OpenStudy (schrodingers_cat):

Yep

OpenStudy (schrodingers_cat):

Your process is right just wrong numbers.

OpenStudy (precal):

so is part b 0?

OpenStudy (precal):

\[\frac{ dA }{ dt}=\frac{ 1 }{ 2 }h \frac{ db }{ dt }+\frac{ 1 }{ 2 }b \frac{ dh}{ dt }\]

OpenStudy (schrodingers_cat):

2a' = yx' + xy' a' = (24(2) + 7(-14/24))/2 = 527/49 ft^2/sec

OpenStudy (precal):

\[\frac{ dA }{ dt }=\frac{ 1 }{ 2 }(24)\frac{ -7 }{ 12 }+\frac{ 1 }{ 2 }(7)(2) \]

OpenStudy (schrodingers_cat):

opps that should be 24 in the denom :P

OpenStudy (schrodingers_cat):

instead of 49

OpenStudy (precal):

24 times 2?

OpenStudy (precal):

sorry I have to relabel my triangle using h and b

OpenStudy (precal):

I believe you are correct

OpenStudy (schrodingers_cat):

Then for c \[\sin(\theta) = x/25\] \[\cos(\theta)\theta' = x'/25\]

OpenStudy (schrodingers_cat):

\[\theta' = 2/(25\cos(.284)) \]

OpenStudy (precal):

ok you lost me with cos how did you get .284? I understood all of the rest. I am using dt notation but I can follow along with your notation

OpenStudy (schrodingers_cat):

\[\sin(\theta) = (7/25)\]

OpenStudy (schrodingers_cat):

\[\theta = \sin^-1(7/25) =.284\]

OpenStudy (precal):

does it matter on part c that it was changed to 9 feet from the wall? Let me rewrite part c

OpenStudy (precal):

Part C. Find the rate at which the angle formed by the ladder and the wall of the house is changing when the base of the ladder is 9 feet from the wall

OpenStudy (precal):

I know the ladder is 25

OpenStudy (schrodingers_cat):

Oh yeah it would change everything I did not see that.

OpenStudy (precal):

ok thanks I think I got it from here. I will post my final solution so you can double check my answer since I tend to mess these up

OpenStudy (schrodingers_cat):

Ok no problem :D

OpenStudy (schrodingers_cat):

Actually you do not need to recalc it would be the same just switch out 7 for 9 when determining the angle then use the equation as x' would remain constant. Sorry I was thinking about how it would affect y :P

OpenStudy (precal):

is dx/dt the given rate of 2 feet per second?

OpenStudy (precal):

that is ok, I have been doing these incorrect because I am making small mistakes

OpenStudy (precal):

ok I got .0857492926

OpenStudy (precal):

for d(theta)/dt

OpenStudy (precal):

rad per second

OpenStudy (schrodingers_cat):

Looks good :)

OpenStudy (precal):

thank you so much I really appreciate this

OpenStudy (schrodingers_cat):

Yep :)

OpenStudy (anonymous):

A solution using Mathematica is attached.